# SpiroLLM: Finetuning Pretrained LLMs to Understand Spirogram Time Series with Clinical Validation in COPD Reporting

### v2

@@ INTRODUCTION
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease characterized by persistent airflow limitation, is one of the leading causes of disability and mortality worldwideVenkatesan2025. In the clinical diagnosis and management of COPD, the Pulmonary Function Test (PFT) and its corresponding spirogram curve are indispensable. They not only represent the gold standard for diagnosis but also serve as a crucial basis for assessing disease severity, monitoring progression, and guiding treatment strategiesagusti2022global. However, the accurate interpretation of spirogram curves and the subsequent drafting of a standardized yet personalized diagnostic report are time-consuming, labor-intensive processes that are highly dependent on the specialized knowledge and long-term experience of clinicians. This reliance on expert resources is particularly pronounced in regions with limited medical access, creating a significant bottleneck in improving the efficiency and standardization of COPD diagnosisstanojevic2022ers.
To address this challenge, researchers have begun exploring the use of Artificial Intelligence (AI) to automate diagnosticsdas2023collaboration. In our prior work, we developed DeepSpiro mei2025deep, a model that demonstrated the feasibility of using deep learning to identify COPD-related features directly from spirogram curves. However, this and other early deep learning models were limited by their "black-box" nature, outputting only simple classification labels. Their inability to provide a rationale for their conclusions has hindered their clinical adoption and trust. More recently, the advent of Large Language Models (LLMs) has shown great promise in addressing this interpretability issue, with their ability to generate logically coherent medical texts that emulate the style of human experts liu2025application. Nevertheless, applying LLMs to generate diagnostic reports directly from raw pulmonary function data still faces three core challenges:
Figure 1: This figure compares three workflows for pulmonary function assessment: the traditional clinical model (A), which relies on cumbersome in-clinic testing; the traditional large language model (B), which cannot understand raw physiological signals; and our proposed SpiroLLM framework (C), which supports at-home self-testing and instant generation of professional reports, significantly improving efficiency.
@@ RESULTS
Figure 2: A schematic diagram of the overall architecture of the SpiroLLM framework. The figure illustrates the complete end-to-end process, from raw pulmonary function test time-series data to the generation of a professional diagnostic report. The blue section represents the SpiroEncoder module, which extracts high-level features from the spirometry curves and performs cross-modal alignment with the large language model via the SpiroProjector. The yellow section is the Query Prompt construction module, which integrates the COPD probability output by the SpiroEncoder, key PFT parameters extracted by SpiroUtils, and the patient’s demographic information to form the model’s input prompt. The green section represents the gold-standard report generation process. This process begins by using the Qwen-VL model to generate morphological descriptions from the pulmonary function curve images, it then incorporates the PFT values extracted by SpiroUtils and introduces relevant domain knowledge through a RAG-based knowledge base system. Finally, all this information is integrated by the DeepSeek V3 model to generate a high-quality, standardized diagnostic report, which serves as the training target output for SpiroLLM.
@@
We evaluated the diagnostic performance of SpiroLLM against the comprehensive set of baselines detailed in the Methods. Table 1 presents the comprehensive diagnostic performance comparison. Statistical significance was assessed using Bootstrap resampling ($n=1000$).
@@
Demographic & Clinical Baselines XGBoost-Demo chen2016xgboost 0.7288 (0.71-0.75) 0.6497 (0.61-0.69) 0.7082 (0.68-0.74) 0.6456 (0.62-0.67) 0.6597 (0.63-0.68)
XGBoost-Anthro chen2016xgboost 0.7368 (0.71-0.76) 0.6816 (0.65-0.71) 0.7703 (0.74-0.80) 0.5953 (0.57-0.62) 0.6783 (0.65-0.70)
XGBoost-Clinical chen2016xgboost 0.8573 (0.84-0.87) 0.8331 (0.81-0.86) 0.7377 (0.71-0.77) 0.8292 (0.81-0.85) 0.7569 (0.73-0.78)
Time-Series Baselines Informer zhou2021informer 0.7687 (0.75-0.79) 0.7098 (0.68-0.74) 0.7351 (0.70-0.77) 0.6800 (0.65-0.71) 0.6898 (0.66-0.71)
PatchTST Yuqietal-2023-PatchTST 0.8073 (0.79-0.83) 0.7810 (0.75-0.81) 0.7389 (0.71-0.77) 0.7378 (0.71-0.76) 0.7161 (0.69-0.74)
TimesNet wu2022timesnet 0.8270 (0.81-0.85) 0.8035 (0.78-0.83) 0.7582 (0.73-0.79) 0.7643 (0.74-0.79) 0.7397 (0.72-0.76)
SpiroEncoder mei2025deep 0.8249 (0.81-0.84) 0.8041 (0.78-0.83) 0.7474 (0.72-0.78) 0.7467 (0.72-0.77) 0.7253 (0.70-0.75)
ResNet18 cosentino2023inference 0.8355 (0.82-0.85) 0.8105 (0.78-0.84) 0.7401 (0.71-0.77) 0.7869 (0.76-0.81) 0.7386 (0.71-0.76)
Multimodal Fusion Baseline DeepSpiro mei2025deep 0.8625 (0.85-0.88) 0.8420 (0.82-0.87) 0.7838 (0.76-0.81) 0.7860 (0.76-0.81) 0.7651 (0.74-0.79)
@@
Figure 3: Interpretable visualization of the SpiroEncoder. Flow-volume curves are color-coded based on attention weights extracted from the final layer of the encoder (red indicates high attention; gray indicates low attention). (a-b) COPD cases: The model exhibits high attention toward the concave descending limb, successfully identifying the characteristic "scooped" pattern associated with airway obstruction. (c-d) Normal cases: Attention remains focused on the descending limb, validating the linear or convex profile associated with normal lung function. This confirms that the model explicitly learns to analyze morphological features indicative of small airway obstruction.
@@
Figure 4: Comparative analysis of SpiroLLM and a baseline model. The figure demonstrates SpiroLLM’s ability to correctly interpret primary diagnostic criteria, while the baseline model is misled by secondary indicators, resulting in an incorrect diagnosis.
@@ METHODS
The application of artificial intelligence in the diagnosis of COPD has made significant progress in recent years. As highlighted in recent reviews by Gompelmann et al. gompelmann2025ai, AI is increasingly transforming COPD management, evolving from simple phenotyping to complex decision support systems. Early research primarily employed traditional machine learning methods, which relied on key metrics extracted from PFTs.
To overcome this limitation, recent studies have shifted their focus to deep learning models capable of learning features directly from raw spirometry time-series data. For instance, in our previous work, we proposed the DeepSpiro model mei2025deep, which processes flow-volume sequences directly to detect COPD. More recently, Cosentino et al. cosentino2023inference validated this paradigm on a large scale using UK Biobank data, demonstrating that Convolutional Neural Networks (CNNs) trained on raw spirograms could identify novel genetic loci. Similarly, Hill et al. hill2025deep introduced Spiro-CLF, a contrastive learning framework that leverages even suboptimal spirometry maneuvers to predict mortality.
Another technical approach involves transforming the time-series waveforms into images for analysis. For example, the AI-PFT-Clin model developed by Jeon et al. jeon2025deep improved the predictive accuracy for COPD acute exacerbation events by fusing clinical variables with images of flow-volume loops. Topole et al. topole2023artificial also converted flow-volume curves into low-resolution images and used a CNN to automatically assess the acceptability of spirometry maneuvers.
@@
Concurrently, the emergence of Multimodal Large Language Models (MLLMs) has provided a powerful technological foundation for the automated generation of complex medical narrative texts liu2025application. These models can integrate and comprehend medical data from diverse sources to produce high-quality reports. In recent surveys, Ye et al. ye2025multimodal and Alsaad et al. alsaad2024multimodal systematically summarized the landscape of MLLMs in healthcare, highlighting their potential to integrate heterogeneous data sources—including Electronic Health Records (EHRs), imaging, and sensor data—to enhance diagnostic precision.
In the field of medical imaging, MLLMs have already been successfully applied to the automated generation of radiology reports, serving as a precedent for other modalities. For instance, the MRG-LLM proposed by Li et al. li2025multimodal can generate targeted and accurate reports for input X-ray images, a technique later extended by Li et al. li2025towards to 3D brain CT scans. These studies demonstrate the robust capability of MLLMs to process complex visual information and translate it into specialized text. Additionally, advanced frameworks like AdaCoMed chen2025multi have explored deeper cross-modal interactions through collaborative learning, while Kaissis et al. kaissis2021end introduced federated learning approaches to safeguard data privacy during multi-center training.
Beyond imaging, significant strides have been made in applying MLLMs to physiological time-series data. Med-PaLM M tu2024towards demonstrated the feasibility of a generalist biomedical AI capable of handling diverse modalities, while HeLM (Belyaeva et al.) belyaeva2023multimodal established a benchmark for grounding LLMs in individual-specific health records.
Architecturally, our SpiroLLM shares structural parallels with these domain-specific MLLMs, employing a Connectivist paradigm—comprising a pre-trained modality encoder and an alignment projector—to bridge the gap between continuous biological signals and discrete textual reasoning. A pertinent example is the GEM model lan2025gemempoweringmllmgrounded, which similarly utilizes a specialized encoder to align ECG signals for clinical interpretation.
However, critical architectural distinctions exist between these diagnostic MLLMs and general time-series adapters. Frameworks like GPT4TS by Zhou et al. zhou2023one and Time-LLM jin2023time have successfully repurposed frozen LLMs for universal time-series analysis. Yet, these models typically utilize a patching mechanism to treat time-series segments as input tokens, optimizing the model primarily for forecasting or anomaly detection tasks. In contrast, they lack the domain-specific alignment required to interpret the global morphological semantics of clinical waveforms for diagnostic reasoning, a critical gap our SpiroLLM aims to fill.
@@
To extract deep feature embeddings $E_{i}$ from the pulmonary function time-series $s_{i}$, this study adopts the DeepSpiro model proposed in prior work mei2025deep as the core time-series encoder. This encoder, denoted as $E_{s}$ in the formula, employs a hybrid CNN-BiLSTM architecture. It first captures key local patterns in the sequence using a one-dimensional Convolutional Neural Network (1D-CNN) and subsequently models the temporal context of these local features using a Bidirectional Long Short-Term Memory (BiLSTM) network.
@@
This step aims to obtain a qualitative description of the patient’s respiratory curve morphology. We first visualize the raw Flow-Volume time-series data to generate standard Flow-Volume curve images. Subsequently, we utilize a powerful multimodal large language model (Qwen2.5-VL-72Bbai2025qwen2), in conjunction with meticulously designed prompts aimed at guiding the model to focus on key morphological features of the curve (e.g., peak shape, degree of concavity in the expiratory limb, etc.), to automatically generate an accurate and objective textual description (see Appendix A for details). This description forms the foundation of the gold-standard report’s section on the patient’s individualized physiological presentation.
@@
To obtain quantitative clinical evidence, we developed a pulmonary function metric calculation tool named SpiroUtils. This tool directly processes the raw time-series data to precisely calculate a series of key pulmonary function parameters, including Forced Vital Capacity (FVC), Forced Expiratory Volume in the first second (FEV1), Forced Expiratory Flow between 25% and 75% of FVC (FEF25%-75%), and FEF75%. More importantly, SpiroUtils integrates the patient’s demographic information (age, gender, height) to calculate the Predicted Value and Z-score for these metrics based on the multi-ethnic reference equations published by the Global Lung Function Initiative in 2012quanjer2012multi.
@@
To ensure that the generated reports comply with the latest clinical guidelines, we constructed a domain knowledge base. The content of this knowledge base is derived from the GOLD 2025 ReportVenkatesan2025. During the report generation process, we employ Retrieval-Augmented Generation (RAG) techniques. Specifically, the morphological descriptions and quantitative metrics from the previous two steps are used as a composite query to retrieve the most relevant knowledge snippets—such as diagnostic criteria, severity grading, and treatment recommendations—from the knowledge base that correspond to the current patient’s condition.
@@
After obtaining the qualitative morphological descriptions of the respiratory curve, the precise quantitative physiological metrics, and the authoritative knowledge from the GOLD standards, we integrate these three components: the morphological description text, the quantitative metrics including Z-scores, and the retrieved relevant knowledge snippets. This integrated information is then formatted according to a meticulously designed structured prompt template (see Appendix B for details). Subsequently, this structured, comprehensive prompt is input into the DeepSeek-V3 modelliu2024deepseek to finally generate a gold-standard diagnostic report $r_{i}^{*}$ that is comprehensive in content, reliable in its conclusions, and aligns with the linguistic style of clinicians. This high-quality report serves as the ground-truth label for the supervised fine-tuning of the main model, thereby ensuring that the model learns accurate diagnostic logic and professional expression.
@@
This study leverages data from 453,558 participants in the UK Biobank (UKB). To ensure the physiological validity and reliability of subsequent analyses and to minimize potential confounding factors, a systematic quality control and sample screening process was implemented. The detailed screening workflow is presented in Appendix Figure E.
@@
This study adopts the parameter-efficient LoRA (Low-Rank Adaptation) strategy to fine-tune the modelhu2022lora. During training, the weights of both the SpiroEncoder and the LLM backbone are kept frozen, where only the parameters of the SpiroProjector and the LoRA adapters in the LLM are updated. The AdamW optimizer is used, with different learning rates set for the SpiroProjector and the LoRA modules. The overall learning rate schedule employs a cosine annealing strategy with a warm-up period, and the optimization objective is the standard language model loss function. To enhance efficiency, bfloat16 mixed-precision is utilized throughout the training process, and an early stopping mechanism is configured to prevent overfitting. The training was conducted on 4 NVIDIA RTX 4090 GPUs.
@@
Time-Series Baselines compared the proposed approach against state-of-the-art waveform encoders, including the flow-volume ResNet18 proposed by Cosentino et alcosentino2023inference., advanced time-series models (Informerzhou2021informer, PatchTSTYuqietal-2023-PatchTST, and TimesNetwu2022timesnet), and SpiroEncoder, the feature extraction backbone employed in the previously developed DeepSpiro frameworkmei2025deep. These models rely exclusively on raw flow-volume curve data.
The Multimodal Fusion Baseline featured DeepSpiromei2025deep, which combines raw flow-volume curves with basic demographics and the FEV1/FVC ratio. This model serves as a representative baseline for traditional feature concatenation strategies, facilitating a rigorous evaluation of the performance gains achieved by the proposed method.
@@
Data from the UK Biobank, which is available after the approval of an application at https://www.ukbiobank.ac.uk. UK Biobank received ethical approval from the National Information Governance Board for Health and Social Care and the National Health Service North West Centre for Research Ethics Committee (Ref: 21/NW/0157).
Code availability
Our SpiroLLM is publicly available at https://github.com/yudaleng/SpiroLLM.

### v3

# arXiv:2507.16145v3[cs.AI] 02 Mar 2026

@@ ABSTRACT
AUTHOR SUMMARY
Chronic Obstructive Pulmonary Disease (COPD) is a leading cause of death and disability worldwide, yet diagnosing it remains a significant challenge. Accurate diagnosis typically requires highly trained specialists to interpret complex charts from breathing tests—a process that is time-consuming and often unavailable in areas with limited medical resources. While artificial intelligence has shown promise in medicine, most existing tools only provide simple "yes or no" classifications without explaining the reasoning behind the diagnosis, which limits their usefulness and trustworthiness for doctors. In this study, we developed a new system called SpiroLLM to bridge this gap. Unlike traditional tools, our approach uses advanced technology to "look" at the shape of a patient’s breathing curves and combine this visual information with standard test numbers. The system then automatically writes a detailed, easy-to-understand clinical report, much like a human expert would. We tested our model using a large dataset from the UK Biobank and found it to be highly accurate and reliable, even when some patient data was missing. This work represents a step forward in digital health, offering a transparent assistant for clinicians that could make expert-level COPD diagnosis more accessible and consistent globally.
@@ INTRODUCTION
Chronic Obstructive Pulmonary Disease (COPD), a major chronic respiratory disease characterized by persistent airflow limitation, is one of the leading causes of disability and mortality worldwide27. In the clinical diagnosis and management of COPD, the Pulmonary Function Test (PFT) and its corresponding spirogram curve are indispensable. They not only represent the gold standard for diagnosis but also serve as a crucial basis for assessing disease severity, monitoring progression, and guiding treatment strategies1. However, the accurate interpretation of spirogram curves and the subsequent drafting of a standardized yet personalized diagnostic report are time-consuming, labor-intensive processes that are highly dependent on the specialized knowledge and long-term experience of clinicians. This reliance on expert resources is particularly pronounced in regions with limited medical access, creating a significant bottleneck in improving the efficiency and standardization of COPD diagnosis24.
To address this challenge, researchers have begun exploring the use of Artificial Intelligence (AI) to automate diagnostics8. In our prior work, we developed DeepSpiro 21, a model that demonstrated the feasibility of using deep learning to identify COPD-related features directly from spirogram curves. However, this and other early deep learning models were limited by their "black-box" nature, outputting only simple classification labels. Their inability to provide a rationale for their conclusions has hindered their clinical adoption and trust. More recently, the advent of Large Language Models (LLMs) has shown great promise in addressing this interpretability issue, with their ability to generate logically coherent medical texts that emulate the style of human experts 20. Nevertheless, applying LLMs to generate diagnostic reports directly from raw pulmonary function data still faces three core challenges:
Fig 1: This figure compares three workflows for pulmonary function assessment: the traditional clinical model (A), which relies on cumbersome in-clinic testing; the traditional large language model (B), which cannot understand raw physiological signals; and our proposed SpiroLLM framework (C), which supports at-home self-testing and instant generation of professional reports, significantly improving efficiency.
@@ RESULTS
Fig 2: A schematic diagram of the overall architecture of the SpiroLLM framework. The figure illustrates the complete end-to-end process, from raw pulmonary function test time-series data to the generation of a professional diagnostic report. The blue section represents the SpiroEncoder module, which extracts high-level features from the spirometry curves and performs cross-modal alignment with the large language model via the SpiroProjector. The yellow section is the Query Prompt construction module, which integrates the COPD probability output by the SpiroEncoder, key PFT parameters extracted by SpiroUtils, and the patient’s demographic information to form the model’s input prompt. The green section represents the gold-standard report generation process. This process begins by using the Qwen-VL model to generate morphological descriptions from the pulmonary function curve images, it then incorporates the PFT values extracted by SpiroUtils and introduces relevant domain knowledge through a RAG-based knowledge base system. Finally, all this information is integrated by the DeepSeek V3 model to generate a high-quality, standardized diagnostic report, which serves as the training target output for SpiroLLM.
@@
We evaluated the diagnostic performance of SpiroLLM against the comprehensive set of baselines detailed in the Methods. Table 1 presents the comprehensive diagnostic performance comparison. Statistical significance was assessed using Bootstrap resampling ($n=10000$).
@@
Demographic & Clinical Baselines XGBoost-Demo 5 0.7288 (0.71-0.75) 0.6497 (0.61-0.69) 0.7082 (0.68-0.74) 0.6456 (0.62-0.67) 0.6597 (0.63-0.68)
XGBoost-Anthro 5 0.7368 (0.71-0.76) 0.6816 (0.65-0.71) 0.7703 (0.74-0.80) 0.5953 (0.57-0.62) 0.6783 (0.65-0.70)
XGBoost-Clinical 5 0.8573 (0.84-0.87) 0.8331 (0.81-0.86) 0.7377 (0.71-0.77) 0.8292 (0.81-0.85) 0.7569 (0.73-0.78)
Time-Series Baselines Informer 30 0.7687 (0.75-0.79) 0.7098 (0.68-0.74) 0.7351 (0.70-0.77) 0.6800 (0.65-0.71) 0.6898 (0.66-0.71)
PatchTST 22 0.8073 (0.79-0.83) 0.7810 (0.75-0.81) 0.7389 (0.71-0.77) 0.7378 (0.71-0.76) 0.7161 (0.69-0.74)
TimesNet 28 0.8270 (0.81-0.85) 0.8035 (0.78-0.83) 0.7582 (0.73-0.79) 0.7643 (0.74-0.79) 0.7397 (0.72-0.76)
SpiroEncoder 21 0.8249 (0.81-0.84) 0.8041 (0.78-0.83) 0.7474 (0.72-0.78) 0.7467 (0.72-0.77) 0.7253 (0.70-0.75)
ResNet18 7 0.8355 (0.82-0.85) 0.8105 (0.78-0.84) 0.7401 (0.71-0.77) 0.7869 (0.76-0.81) 0.7386 (0.71-0.76)
Multimodal Fusion Baseline DeepSpiro 21 0.8625 (0.85-0.88) 0.8420 (0.82-0.87) 0.7838 (0.76-0.81) 0.7860 (0.76-0.81) 0.7651 (0.74-0.79)
@@
Fig 3: Interpretable visualization of the SpiroEncoder. Flow-volume curves are color-coded based on attention weights extracted from the final layer of the encoder (red indicates high attention; gray indicates low attention). (a-b) COPD cases: The model exhibits high attention toward the concave descending limb, successfully identifying the characteristic "scooped" pattern associated with airway obstruction. (c-d) Normal cases: Attention remains focused on the descending limb, validating the linear or convex profile associated with normal lung function. This confirms that the model explicitly learns to analyze morphological features indicative of small airway obstruction.
@@
Fig 4: Comparative analysis of SpiroLLM and a baseline model. The figure demonstrates SpiroLLM’s ability to correctly interpret primary diagnostic criteria, while the baseline model is misled by secondary indicators, resulting in an incorrect diagnosis.
@@ METHODS
The application of artificial intelligence in the diagnosis of COPD has made significant progress in recent years. As highlighted in recent reviews by Gompelmann et al. 9, AI is increasingly transforming COPD management, evolving from simple phenotyping to complex decision support systems. Early research primarily employed traditional machine learning methods, which relied on key metrics extracted from PFTs.
To overcome this limitation, recent studies have shifted their focus to deep learning models capable of learning features directly from raw spirometry time-series data. For instance, in our previous work, we proposed the DeepSpiro model 21, which processes flow-volume sequences directly to detect COPD. More recently, Cosentino et al. 7 validated this paradigm on a large scale using UK Biobank data, demonstrating that Convolutional Neural Networks (CNNs) trained on raw spirograms could identify novel genetic loci. Similarly, Hill et al. 11 introduced Spiro-CLF, a contrastive learning framework that leverages even suboptimal spirometry maneuvers to predict mortality.
Another technical approach involves transforming the time-series waveforms into images for analysis. For example, the AI-PFT-Clin model developed by Jeon et al. 13 improved the predictive accuracy for COPD acute exacerbation events by fusing clinical variables with images of flow-volume loops. Topole et al. 25 also converted flow-volume curves into low-resolution images and used a CNN to automatically assess the acceptability of spirometry maneuvers.
@@
Concurrently, the emergence of Multimodal Large Language Models (MLLMs) has provided a powerful technological foundation for the automated generation of complex medical narrative texts 20. These models can integrate and comprehend medical data from diverse sources to produce high-quality reports. In recent surveys, Ye et al. 29 and Alsaad et al. 2 systematically summarized the landscape of MLLMs in healthcare, highlighting their potential to integrate heterogeneous data sources—including Electronic Health Records (EHRs), imaging, and sensor data—to enhance diagnostic precision.
In the field of medical imaging, MLLMs have already been successfully applied to the automated generation of radiology reports, serving as a precedent for other modalities. For instance, the MRG-LLM proposed by Li et al. 18 can generate targeted and accurate reports for input X-ray images, a technique later extended by Li et al. 17 to 3D brain CT scans. These studies demonstrate the robust capability of MLLMs to process complex visual information and translate it into specialized text. Additionally, advanced frameworks like AdaCoMed 6 have explored deeper cross-modal interactions through collaborative learning, while Kaissis et al. 15 introduced federated learning approaches to safeguard data privacy during multi-center training.
Beyond imaging, significant strides have been made in applying MLLMs to physiological time-series data. Med-PaLM M 26 demonstrated the feasibility of a generalist biomedical AI capable of handling diverse modalities, while HeLM (Belyaeva et al.) 4 established a benchmark for grounding LLMs in individual-specific health records.
Architecturally, our SpiroLLM shares structural parallels with these domain-specific MLLMs, employing a Connectivist paradigm—comprising a pre-trained modality encoder and an alignment projector—to bridge the gap between continuous biological signals and discrete textual reasoning. A pertinent example is the GEM model 16, which similarly utilizes a specialized encoder to align ECG signals for clinical interpretation.
However, critical architectural distinctions exist between these diagnostic MLLMs and general time-series adapters. Frameworks like GPT4TS by Zhou et al. 31 and Time-LLM 14 have successfully repurposed frozen LLMs for universal time-series analysis. Yet, these models typically utilize a patching mechanism to treat time-series segments as input tokens, optimizing the model primarily for forecasting or anomaly detection tasks. In contrast, they lack the domain-specific alignment required to interpret the global morphological semantics of clinical waveforms for diagnostic reasoning, a critical gap our SpiroLLM aims to fill.
@@
To extract deep feature embeddings $E_{i}$ from the pulmonary function time-series $s_{i}$, this study adopts the DeepSpiro model proposed in prior work 21 as the core time-series encoder. This encoder, denoted as $E_{s}$ in the formula, employs a hybrid CNN-BiLSTM architecture. It first captures key local patterns in the sequence using a one-dimensional Convolutional Neural Network (1D-CNN) and subsequently models the temporal context of these local features using a Bidirectional Long Short-Term Memory (BiLSTM) network.
@@
This step aims to obtain a qualitative description of the patient’s respiratory curve morphology. We first visualize the raw Flow-Volume time-series data to generate standard Flow-Volume curve images. Subsequently, we utilize a powerful multimodal large language model (Qwen2.5-VL-72B3), in conjunction with meticulously designed prompts aimed at guiding the model to focus on key morphological features of the curve (e.g., peak shape, degree of concavity in the expiratory limb, etc.), to automatically generate an accurate and objective textual description (see Appendix A for details). This description forms the foundation of the gold-standard report’s section on the patient’s individualized physiological presentation.
@@
To obtain quantitative clinical evidence, we developed a pulmonary function metric calculation tool named SpiroUtils. This tool directly processes the raw time-series data to precisely calculate a series of key pulmonary function parameters, including Forced Vital Capacity (FVC), Forced Expiratory Volume in the first second (FEV1), Forced Expiratory Flow between 25% and 75% of FVC (FEF25%-75%), and FEF75%. More importantly, SpiroUtils integrates the patient’s demographic information (age, gender, height) to calculate the Predicted Value and Z-score for these metrics based on the multi-ethnic reference equations published by the Global Lung Function Initiative in 201223.
@@
To ensure that the generated reports comply with the latest clinical guidelines, we constructed a domain knowledge base. The content of this knowledge base is derived from the GOLD 2025 Report27. During the report generation process, we employ Retrieval-Augmented Generation (RAG) techniques. Specifically, the morphological descriptions and quantitative metrics from the previous two steps are used as a composite query to retrieve the most relevant knowledge snippets—such as diagnostic criteria, severity grading, and treatment recommendations—from the knowledge base that correspond to the current patient’s condition.
@@
After obtaining the qualitative morphological descriptions of the respiratory curve, the precise quantitative physiological metrics, and the authoritative knowledge from the GOLD standards, we integrate these three components: the morphological description text, the quantitative metrics including Z-scores, and the retrieved relevant knowledge snippets. This integrated information is then formatted according to a meticulously designed structured prompt template (see Appendix B for details). Subsequently, this structured, comprehensive prompt is input into the DeepSeek-V3 model19 to finally generate a gold-standard diagnostic report $r_{i}^{*}$ that is comprehensive in content, reliable in its conclusions, and aligns with the linguistic style of clinicians. This high-quality report serves as the ground-truth label for the supervised fine-tuning of the main model, thereby ensuring that the model learns accurate diagnostic logic and professional expression.
@@
This study leverages data from 453,558 participants in the UK Biobank (UKB). To ensure the physiological validity and reliability of subsequent analyses and to minimize potential confounding factors, a systematic quality control and sample screening process was implemented. The detailed screening workflow is presented in Appendix E Figure.
@@
This study adopts the parameter-efficient LoRA (Low-Rank Adaptation) strategy to fine-tune the model12. During training, the weights of both the SpiroEncoder and the LLM backbone are kept frozen, where only the parameters of the SpiroProjector and the LoRA adapters in the LLM are updated. The AdamW optimizer is used, with different learning rates set for the SpiroProjector and the LoRA modules. The overall learning rate schedule employs a cosine annealing strategy with a warm-up period, and the optimization objective is the standard language model loss function. To enhance efficiency, bfloat16 mixed-precision is utilized throughout the training process, and an early stopping mechanism is configured to prevent overfitting. The training was conducted on 4 NVIDIA RTX 4090 GPUs.
@@
Time-Series Baselines compared the proposed approach against state-of-the-art waveform encoders, including the flow-volume ResNet18 proposed by Cosentino et al7., advanced time-series models (Informer30, PatchTST22, and TimesNet28), and SpiroEncoder, the feature extraction backbone employed in the previously developed DeepSpiro framework21. These models rely exclusively on raw flow-volume curve data.
The Multimodal Fusion Baseline featured DeepSpiro21, which combines raw flow-volume curves with basic demographics and the FEV1/FVC ratio. This model serves as a representative baseline for traditional feature concatenation strategies, facilitating a rigorous evaluation of the performance gains achieved by the proposed method.
@@
This research has been conducted using the UK Biobank Resource under Application Number 90018. Researchers can apply for data access via the UK Biobank Access Management System: https://www.ukbiobank.ac.uk. The authors confirm that they did not have any special access privileges that others would not have. UK Biobank received ethical approval from the National Information Governance Board for Health and Social Care and the National Health Service North West Centre for Research Ethics Committee (Ref: 21/NW/0157). The analytic code used in this study is fully available at: https://github.com/yudaleng/SpiroLLM.